OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (anonymous):

no solution in this equation bcz x get lost in this equation

OpenStudy (karatechopper):

−2(x + 3) = −2(x + 1) − 4 Distributive property −2(x + 3) = −2(x + 1) − 4 -2x-6=-2x-2-4 Add/subtract like terms on their respective sides -2x-6=-2x-2-4 -2x-6=-2x-6 Add 6 to both sides -2x-6=-2x-6 +6 +6 -------------- -2x=-2x Divide by -2 on both sides -2x=-2x -2x/-2=-2x/-2 x=x All real numbers Thanks :) @yummydum

OpenStudy (anonymous):

there ya go ;) @karatechopper

hero (hero):

Good strategy @karatechopper ...ending with x = x. I'll have to use that.

OpenStudy (karatechopper):

I would medal that too^

hero (hero):

lol actually, I made a mistake hope nobody sees it

hero (hero):

−2(x + 3) = −2(x + 1) − 4 Divide both sides by -2: x + 3 = x + 1 + 2 x + 3 = x + 3 x - x = 3 - 3 0 = 0